4a+7b=22
8a+3c=7
6b+3c+3

Answer by checkley75(3666)

4a+7b=22
8a+3c=7
-(6b+3c=3) I'll assume the + should have been an = sign. Now subtract.
--------
8a-6b=4 Now multiply the first equation by -2 & add.
-8a-14b=-44
--------------
-20b=-40
b=-40/-20
b=2 ans.
8a-6*2=4
8a-12=4
8a=4+12
8a=16
a=16/8
a=2 ans.
6*2+3c=3
12+3c=3
3c=3-12
3c=-9
c=-9/3
c=-3 ans.